OpenStudy (anonymous):

4. The area of a garden is given by the trinomial g2 – 2g – 24. The garden’s length is g + 4. What is the garden’s width? (1 point)

OpenStudy (pfenn1):

A\[\text{Area }=\text{ width } \times \text { length }\]

OpenStudy (pfenn1):

\[g^2-2g-24=(g+4)(g+a)\]where a must fit the following :\[4+a=-2\]\[4a=-24\]